A patient arrives at a medical center and has to make decisions about their treatment schedule. Upon arrival, the patient has to choose whether they will go to the waiting room for their regular provider, where the expected wait time is 45 minutes, or to an overflow waiting room where the expected wait time is 10 minutes. If they go to the waiting room for their regular provider, when they are called in they have to choose if they will see an MD or an NP. It is expected that it will take an MD 30 minutes to enter the exam room, and an NP will enter after 10 minutes. The appointment duration with an MD averages 15 minutes, and for an NP the average time is 25 minutes. After an appointment with either an MD or an NP, the patient goes to Billing Station 1 where it takes them 10 minutes to checkout, after which they can go home. If the patient first chooses to go to the overflow waiting room, when they are called in to see a provider they will meet with a PRN, who will enter the exam room after 15 minutes. After being screened by the PRN, the patient must choose if they want to see an MD, an NP, or a PA. It is expected that it will take 60 minutes for an MD to enter the exam room, 15 minutes for an NP, and 40 minutes for a PA. The average exam times for an MD or NP are the same as above, and the average exam time for an NP is 20 minutes. If the patient has an appointment with a PA, they can choose to go to Billing Station 1, or they can go to Billing Station 2 where it takes on average 25 minutes to check out. After check-out, the patient can go home. Formulate a network model that will allow the patient to determine the sequence of decisions they should make to minimize their total time in the medical center. [Hint: it can be modeled a shortest path problem.]
Answer #1

Network Model

A network model is a database model that is designed as a flexible approach to representing objects and their relationships.

Below image is the graphical explanation of the above scenario using network model

NETWORK MODEL I Patient 45min lomin waiting room Oreatlow room! 15 min 30min/ zomin . INP [PRN PRN L I 15 20 THO Wp [PAL 15 mNow we will calculate the total time with help of network model ollagram. Note - waiting room denoted as wR and Overflow roomOverfull room - PNR = 10+15+15 NP Bluing 1 +25+10 = 1 hour 15 min Overfull room PNR PA y Biling 1 = 10 +15+40+20+ 10 - = 95mi